# references.bib

### v2

@@ 1 Introduction
Atmospheric aerosols affect climate directly by scattering and absorbing light [chylek1995effect] and indirectly by enhancing cloud formation [lohmann2005global]. One of the components of atmospheric aerosols and a strong light absorber is soot [haywood2000estimates]. Soot is formed upon combustion of carbon-containing fuels in a multi-step process involving thermolysis of the fuel, PAH formation, nucleation, and oxidation [thomson2018radical, johansson2018resonance]. As a result, fractal-like structures consisting mostly of elemental carbon are produced [demirdjian2007heterogeneities]. The micro-structure of carbonaceous phase can be either amorphous, fullerenic, or a combination of both, depending on conditions, under which the aggregate was formed [grieco2000fullerenic].
@@ 2.1.1 Development of the contact model
In a system of $N$ particles we can have up to $\frac{1}{2}N(N-1)$ binary interactions. Let us pick two particles $i$ and $j$ from system and consider their common reference frame. There are four types of relative motion that particles $i$ and $j$ can undergo, corresponding to the four degrees of freedom (DOFs) that we would like to control (Figure 2). These DOFs are (a) normal translation, (b) torsion, (c) tangential translation, and (d) rolling.
@@
In order to control each DOF listed in Figure 2, we insert four springs between particles $i$ and $j$. The rates of stretching/contraction of these four (normal, torsional, tangential, and rolling) springs are given by $\mathbf{v}_{ij,\rm n}$, $r\boldsymbol{\upomega}_{ij,\rm n}$, $\mathbf{v}_{ij,\rm t}$, and $\boldsymbol{\upomega}_{ij,\rm r}\times a\mathbf{n}$ respectively. Length of the normal spring can be computed directly at every time step from positions of particles $i$ and $j$. As to the remaining three springs, their lengths start at zero when the contact is established and need to be incremented throughout the duration of the contact. Let us define a generic term for a vector representing one of the three springs (torsional, tangential, or rolling), $𝛏\boldsymbol{\upxi}$, and its rate of change, $𝛏\dot{\boldsymbol{\upxi}}$. Then the algorithm at every time step is as follows:
@@ 2.1.2 Elastic bond model
Soot restructuring is limited by the strength of elastic bonds between primary particles [weber1997situ]. It is known that most-to-all neighboring primary particles in a soot aggregate are initially bonded with rigid necks [rothenbacher2008fragmentation], but some of the bonds break under the pressure applied by the AFM tip or the stress imposed by the surface tension of a condensed liquid. In lieu of a bond strength parametrization, we prescribe a certain fraction of broken bonds at the start of every simulation. Then, aggregate fragments can fold, slide, and roll relative to each other only at the non-bonded contacts. If the fraction of broken bonds is zero, then all adjacent primary particles in the aggregate are bonded with necks and no folding, sliding, or rolling is possible.
@@ 3.1.1 Aggregate deposition
This simulation stage begins with an aggregate, where all neighboring primary particles are bonded with necks. The substrate is introduced into the simulation as a static plane. It can exert forces on the granular system, like van der Waals attraction, normal force, and friction, but cannot accelerate itself. AFM tip is not present at this simulation stage. The substrate plane is normal to the z-axis and is located at $z=0$. A numerically generated fractal aggregate is loaded and positioned such that the lowest monomer (with respect to the z-axis) is only slightly above the substrate plane (Figure 4a). Each primary particle is initialized with $1\ \rm m/s$ velocity in the negative-z direction.
@@ 3.1.2 Aggregate indentation
The velocity profile has a sigmoid shape. The tip moves with a constant velocity of $1.5\ \rm m/s$ towards the substrate, then at time $t=10^{-8}\ \rm s$ it smoothly reverses its motion and starts moving away from the substrate with a constant velocity of $1.5\ \rm m/s$. Due to strong adhesion, a branch of the aggregate attaches to and follows the tip (Figure 5b). The force exerted by the AFM tip on the granular system is recorded as a function of tip height.
@@
Figure 6: (a) A simulated force-displacement curve (with the transfer function filter applied) and (b) the simulated force-displacement curve with a reduced number of points plotted next to an example experimental force-displacement curve
The simulated force-displacement curve has the expected sawtooth profile. The magnitude of force peaks, their number, and the displacement at which the aggregate is completely severed are comparable between experimental and simulated data. The outcome of indentation simulations is mostly affected by (1) initial morphology of the aggregate and (2) the Hamaker constant used to model van der Waals attraction between monomers. Aggregate morphology controls the number of peaks seen on a force-displacement curve, as the probability of multiple collisions between branches occurring is higher in a larger and more lacey aggregate. Hamaker constant controls the magnitude of the force peaks on the force-displacement curve. The good agreement between simulations and experiments can be explained by us using realistic fractal parameters (Section 2.1.6) and a Hamaker constant value for graphite consistent with literature (Section 2.1.4).
@@ 3.2 Restructuring simulation
where $c$ is the average coordination number of the aggregate and $N_{\mathrm{nb},i}$ is the number of neighbors that particle $i$ has. For the purpose of finding $N_{\mathrm{nb},i}$, we consider as neighbors two particles, which are separated by less that $1\%$ of primary particle diameter. In restructuring simulations, 50 different aggregates with the same size (150 primary particles) and fractal parameters ($1.8$ fractal dimension and $1.4$ pre-exponential factor) were restructured with the same coating force for the same duration. Average coordination numbers of aggregates with various necking fractions are presented in Figure 8a,b. Coordination number is descriptive of local morphological changes in an aggregate (inset in Figure 8b). We also compute convexity of the restructuring aggregates, which in three dimensions is defined as the ratio of the volume of an aggregate to the volume of the convex hull constructed around the aggregate [cgal:hs-ch3-24a]. Convexity is descriptive of global morphological changes in the aggregate (inset in Figure 8d). Convexity data is presented in Figure 8c,d.

### v3

@@ 1 Introduction
Atmospheric aerosols affect climate directly by scattering and absorbing light [chylek1995effect] and indirectly by enhancing cloud formation [lohmann2005global]. One of the components of atmospheric aerosols and a strong light absorber is soot [haywood2000estimates]. Soot is formed upon combustion of carbon-containing fuels in a multi-step process involving thermolysis of the fuel, PAH formation, nucleation, and oxidation [thomson2018radical, johansson2018resonance]. As a result, fractal-like structures consisting mostly of elemental carbon are produced [demirdjian2007heterogeneities]. The micro-structure of carbonaceous phase can be either amorphous, fullerenic, or a combination of both, depending on conditions under which the aggregate was formed [grieco2000fullerenic].
@@ 2.1.1 Development of the contact model
In a system of $N$ particles we can have up to $\frac{1}{2}N(N-1)$ binary interactions. Let us pick two particles $i$ and $j$ from system and consider their common reference frame. There are four types of relative motion that particles $i$ and $j$ can undergo, corresponding to the four degrees of freedom that we would like to control (Figure 2). These degrees of freedom are (a) normal translation, (b) torsion, (c) tangential translation, and (d) rolling.
@@
In order to control each degree of freedom listed in Figure 2, we insert four springs between particles $i$ and $j$. The rates of stretching/contraction of these four (normal, torsional, tangential, and rolling) springs are given by $\mathbf{v}_{ij,\rm n}$, $r\boldsymbol{\upomega}_{ij,\rm n}$, $\mathbf{v}_{ij,\rm t}$, and $\boldsymbol{\upomega}_{ij,\rm r}\times a\mathbf{n}$ respectively. Length of the normal spring can be computed directly at every time step from positions of particles $i$ and $j$. As to the remaining three springs, their lengths start at zero when the contact is established and need to be incremented throughout the duration of the contact. Let us define a generic term for a vector representing one of the three springs (torsional, tangential, or rolling), $𝛏\boldsymbol{\upxi}$, and its rate of change, $𝛏\dot{\boldsymbol{\upxi}}$. Then the algorithm at every time step is as follows:
@@ 2.1.2 Elastic bond model
Soot restructuring is limited by the strength of elastic bonds between primary particles [weber1997situ]. It is known that most-to-all neighboring primary particles in a soot aggregate are initially bonded with rigid necks [rothenbacher2008fragmentation], but some of the bonds break under the pressure applied by the AFM tip or the stress imposed by the surface tension of a condensed liquid. Neck fracture parameterization should consider microstructure and possible anisotropy of carbonaceous materials, but that is beyond the scope of this study, which focuses primarily on contact model development. In lieu of a bond strength parametrization, we prescribe a certain fraction of broken bonds at the start of every simulation. Then, aggregate fragments can fold, slide, and roll relative to each other only at the non-bonded contacts. If the fraction of broken bonds is zero, then all adjacent primary particles in the aggregate are bonded with necks and no folding, sliding, or rolling is possible.
@@ 3.1.1 Aggregate deposition
This simulation stage begins with an aggregate, where all neighboring primary particles are bonded with necks. The substrate is introduced into the simulation as a static plane. It can exert forces on the granular system, like van der Waals attraction, normal force, and friction, but cannot accelerate itself. AFM tip is not present at this simulation stage. The substrate plane is normal to the z-axis and is located at $z=0$. A numerically generated fractal aggregate is loaded and positioned such that the lowest monomer (with respect to the z-axis) is only slightly above the substrate plane (Figure 4a). Each primary particle is initialized with $1\ \rm m/s$ velocity in the negative-z direction. This is a reasonable deposition velocity at which no change in aggregate morphology upon impact is expected. e.g., as reported by Rothenbacher et al. [rothenbacher2008fragmentation], fully necked aggregates do not fragment at impact velocities up to $300\rm\ m/s$.
@@ 3.1.2 Aggregate indentation
The velocity profile has a sigmoid shape. The tip moves with a constant velocity of $1.5\ \rm m/s$ towards the substrate, then at time $t=10^{-8}\ \rm s$ it smoothly reverses its motion and starts moving away from the substrate with a constant velocity of $1.5\ \rm m/s$. Due to strong adhesion, a branch of the aggregate attaches to and follows the tip (Figure 5b). The force exerted by the AFM tip on the granular system is recorded as a function of tip height. It is important to note that in AFM force spectroscopy experiments, the probe moves at much lower velocities, on the order of microns per second, and the timescale of experiments is proportionally larger. However, running simulations for such long time periods is not feasible and numerical studies resort to stretching velocities on the order of meters per second [laube2018new].
@@
Figure 6: A simulated force-displacement curve (with the transfer function filter applied) plotted next to an example experimental force-displacement curve
The simulated force-displacement curve has the expected sawtooth profile. The magnitude of force peaks, their number, and the displacement at which the aggregate is completely severed are comparable between experimental and simulated data. The outcome of indentation simulations is mostly affected by (1) initial morphology of the aggregate and (2) the Hamaker constant used to model van der Waals attraction between monomers. Aggregate morphology controls the number of peaks seen on a force-displacement curve, as the probability of multiple collisions between branches occurring is higher in a larger and more lacey aggregate. Hamaker constant controls the magnitude of the force peaks on the force-displacement curve. The good agreement between simulations and experiments can be explained by us using realistic fractal parameters (Section 2.1.6), a Hamaker constant value for graphite consistent with literature (Section 2.1.4), and aggregates with approximately the same number of primary particles as in our experiments. Varying the number of monomers in the aggregate may affect the total number of sawtooth peaks on the force-displacement curve.
@@ 3.2 Restructuring simulation
where $c$ is the average coordination number of the aggregate and $N_{\mathrm{nb},i}$ is the number of neighbors that particle $i$ has. For the purpose of finding $N_{\mathrm{nb},i}$, we consider as neighbors two particles, which are separated by less that $1\%$ of primary particle diameter. In restructuring simulations, 50 different aggregates with the same size (150 primary particles) and fractal parameters ($1.8$ fractal dimension and $1.4$ pre-exponential factor) were restructured with the same coating force for the same duration. Average coordination numbers of aggregates with various necking fractions are presented in Figure 8a,b. Coordination number is descriptive of local morphological changes in an aggregate (inset in Figure 8b). We also compute convexity of the restructuring aggregates, which in three dimensions is defined as the ratio of the volume of an aggregate to the volume of the convex hull constructed around the aggregate [cgal:hs-ch3-24a]. Convexity is descriptive of global morphological changes in the aggregate (inset in Figure 8d). Convexity data is presented in Figure 8c,d. Three-dimensional convexity presented in Figure 8 cannot be trivially related to experimental convexity derived from images, as in experimental studies convexity is two-dimensional (ratio of aggregate projected area to the area of the convex polygon) and is an inherently different parameter. Terminal two-dimensional convexity can be estimated by averaging convexities of aggregate projections onto the $xy$, $xz$, and $yz$ planes. For aggregates with necking fractions of $0\%$ and $70\%$, terminal 2D convexity was found to be $0.81\pm 0.03$ and $0.80\pm 0.03$ respectively. Computed 2D convexities lie in the range from $0.75$ to $0.87$, reported for restructured aggregates in experimental studies [chen2016unexpected, chen2018single].